OpenStudy (anonymous):

What is the solution set of |x - 6| = 4?

OpenStudy (anonymous):

forgetting the absolute symbol for a moment, what value would you have to make x to make the statement true?

jigglypuff314 (jigglypuff314):

|x - 6| = 4 can be separated into to equations x - 6 = 4 and -(x - 6) = 4 - x + 6 = 4 then solve for x for both and then you have your answers.

OpenStudy (anonymous):

10 10 or -10 -2 or -10 2 or 10

OpenStudy (anonymous):

them are the answers i have

jigglypuff314 (jigglypuff314):

which one of the answers is your choice then?

OpenStudy (anonymous):

The answer is x=10 x-6=4 x+6=+6 x =10 you just trying to get the variable by its self. so always do the opposite of the number that's the closest to the variable. (you closest friend stays to the end) so same thing with the variables!